# Modeling phase transformations in Mn-rich disordered rocksalt cathodes with charge-informed machine-learning interatomic potentials

### v1

@@ V Acknowledgments
This work was primarily supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC0205CH11231 (Materials Project program KC23MP), and the Assistant Secretary for Energy Efficiency and Renewable Energy, Vehicle Technologies Office, under the Advanced Battery Materials Research (BMR) Program of the US Department of Energy (DOE) under contract No. DE-AC0205CH11231. The computational modeling in this work was supported by the computational resources provided by the Extreme Science and Engineering Discovery Environment (XSEDE), supported by National Science Foundation grant number ACI1053575; the National Energy Research Scientific Computing Center (NERSC); and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory. The authors thank Zijian Cai for the valuable discussions.
@@ Appendix A DFT Calculations
All DFT calculations were performed with the VASP package using the projector-augmented wave method [51, 52], a plane-wave basis set with an energy cutoff of 680 eV, and a reciprocal space discretization of 25 k-points per Å-1. The calculations were converged to eV in total energy for electronic loops and 0.02 eV/Å in interatomic forces for ionic loops. The regularized strongly constrained and appropriately normed meta-GGA exchange-correlation functional (r2SCAN) [36, 53] was used with consistent computational settings as MPScanRelaxSet [54].
@@ Appendix C Li-vacancy phase diagram
To construct the simplified intercalation voltage profile calculations, we used an enumeration method for the generation of the delithiated structure and the construction of Li-vacancy phase diagrams. For each delithiation step as an example, it follows:
(1) Select an initial structure with composition LiTMO2, relax the structure using the product-CHGNet, and obtain the internal energy .
@@
For the lithiation process, the selection criterion is changed to find the lithiated structure with the maximum voltage. It is important to note that different 0-TM channels may share the same Li atoms in an FCC framework and therefore the conversion of different 0-TM to Li thus may result in delithiated structures with varying Li concentrations in Step (2). It is more appropriate to use the minimum/maximum voltage as a selection criterion rather than the energy itself for suggesting the next step in the enumeration method.

### v2

@@ V Acknowledgments
This work was primarily supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC0205CH11231 (Materials Project program KC23MP), and the Assistant Secretary for Energy Efficiency and Renewable Energy, Vehicle Technologies Office, under the Advanced Battery Materials Research (BMR) Program of the US Department of Energy (DOE) under contract No. DE-AC0205CH11231. The computational modeling in this work was supported by the computational resources provided by the Extreme Science and Engineering Discovery Environment (XSEDE), supported by National Science Foundation grant number ACI1053575; the National Energy Research Scientific Computing Center (NERSC); the National Renewable Energy Laboratory (NREL) clusters under the drx allocation, and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory. The authors thank Zijian Cai for the valuable discussions.
Data Availability
The supporting data and code are available at https://github.com/zhongpc/drx2delta.
@@ Appendix A DFT Calculations
All DFT calculations were performed with the VASP package using the projector-augmented wave method [57, 58], a plane-wave basis set with an energy cutoff of 680 eV, and a k-point sampling grid spacing of 0.25 Å-1. The calculations were converged to eV in total energy for electronic loops and 0.02 eV/Å in interatomic forces for ionic loops. The regularized strongly constrained and appropriately normed meta-GGA exchange-correlation functional (r2SCAN) [37, 59] was used with consistent computational settings as MPScanRelaxSet [60].
@@ Appendix C Li-vacancy phase diagram
To construct the simplified intercalation voltage profiles, we used an enumeration method for the generation of the delithiated structure and the construction of Li-vacancy phase diagrams. For each delithiation level, the following steps are performed:
(1) Select an initial structure with composition LiTMO2, relax the structure using Product-CHGNet, and obtain the internal energy .
@@
For the lithiation process, the selection criterion is changed to find the lithiated structure with the maximum voltage. It is important to note that different 0-TM sites may share the same Li atoms in an FCC framework and therefore the conversion of different 0-TM to Li thus may result in delithiated structures with varying Li concentrations in Step (2). It is more appropriate to use the minimum/maximum voltage as a selection criterion rather than the energy itself for suggesting the next step in the enumeration method.